Can you explain the steps in-between these steps so I understand why this answer works?
8x + e = (ex + 8)ln(8)
to
8x + e = exln(8) + 8ln(8)
to
8x - exln(8) = 8ln(8) - e
to
x = ( 8ln(8) - e ) / ( 8 - eln(8) )
to
x = 5.93

Answer by ewatrrr(24785)

8x + e = (ex + 8)ln(8)
Removing the Parenthesis by multiplying by ln(8)
8x + e = exln(8) + 8ln(8)
Subtracting e and exln(8) fron both sides of = sign
8x - exln(8) = 8ln(8) - e
Distributive Property
x(8 - eln(8)) = 8ln(8) - e
Dividing both sides by (8-eln(8))
x = ( 8ln(8) - e ) / ( 8 - eln(8) )
Using Calculator: e = 2.718 and ln(8) = 2.079
x = 5.93